# Zeroth-order log-concave sampling

### v1

@@ 4.1 Proximal sampler for truncated Gaussian
Thus, $\mathbb{E}_{\mu_{k}^{Y}}[(1-\ell)^{\tau}]\leq 1/2$, and the expected number of trials during the $k$-th iteration is bounded by $2\tau$ as in §2.2. Combining all these together, we can prove Theorem 1.4 on the query complexity of $\mathsf{PS}_{\textup{Gauss}}$.
@@
For $M:=\mathcal{R}_{2\wedge q_{0}}(\mu_{0}^{X}\mathbin{\|}\mu^{X})$, we set $h^{-1}\asymp d^{2}M$ and $\tau\asymp Me^{M}$. As mentioned in the mixing analysis, $\mathsf{PS}_{\textup{Gauss}}$ returns a sample $X_{N}$ such that $\mathcal{R}_{q}(\operatorname{law}X_{N}\mathbin{\|}\mu^{X})\leq\varepsilon$ by iterating
$N\lesssim Md^{2}\sigma^{2}\log\frac{qM}{(2\wedge q_{0}-1)\,\varepsilon}\quad\text{times}\,.$
Thus, the expected number of total queries during $N$ iterations is
$2\tau N\lesssim M^{2}e^{M}d^{2}\sigma^{2}\log\frac{qM}{(2\wedge q_{0}-1)\,\varepsilon}=\widetilde{O}\bigl{(}d^{2}\sigma^{2}\,e^{\mathcal{R}_{2\wedge q_{0}}(\mu_{0}^{X}\mathbin{\|}\mu^{X})}\log\frac{q}{(2\wedge q_{0}-1)\,\varepsilon}\bigr{)}\,,$
which completes the proof. ∎
@@ 4.2 Functional inequalities to closeness and early-stopping
To avoid any regularity issues, we work with mollifications of log-concave distributions. As detailed in [Eva10, §C.4], the standard mollifier $\eta\in C^{\infty}(\mathbb{R}^{d})$ is defined as
$\eta(x):=\begin{cases}C\exp\bigl{(}\frac{1}{\lVert x\rVert^{2}-1}\bigr{)}&\lVert x\rVert\leq 1\,,\\ 0&\lVert x\rVert>1\,,\end{cases}$ (4.2)
@@ A Rényi version of LSI was established in [VW23, Lemma 5] that for $q>1$ and smooth $\mu/\nu$,
Let $\pi$ be a log-concave probability measure over $\mathbb{R}^{d}$ with $\mathbb{E}_{\pi}[|\cdot|^{2}]=R^{2}$. For $\alpha,\sigma^{2}>0$ and $q>1$, probability measures $\pi\gamma_{\sigma^{2}}$ and $\pi\gamma_{\sigma^{2}(1+\alpha)}$ satisfy
@@
$\frac{\int|\cdot|^{2}\,\gamma_{\delta}\,\mathrm{d}\rho\,}{\int\gamma_{\delta}\,\mathrm{d}\rho}\leq\frac{\int|\cdot|^{2}\,\mathrm{d}\rho}{\int\mathrm{d}\rho}=\mathbb{E}_{\rho}[|\cdot|^{2}]\,.$
Observe that $\operatorname{cov}_{\rho}(f,g)=\frac{1}{2}\,\mathbb{E}_{(X,Y)\sim\rho^{\otimes 2}}[\{f(X)-f(Y)\}\,\{g(X)-g(Y)\}]$. As $f(\cdot)=|\cdot|^{2}$ increases in ||$\cdot$|| ⋅ | while $g=\gamma_{\delta}$ decreases, $0\geq\operatorname{cov}_{\rho}(f,g)=\mathbb{E}_{\rho}[fg]-\mathbb{E}_{\rho}f\,\mathbb{E}_{\rho}g$, which proves the claim. Using this fact with $\rho=\pi$, we complete the proof of the main claim. ∎
@@
Let $\pi$ be a log-concave probability measure over $\mathbb{R}^{d}$ with $\mathbb{E}_{\pi}[|\cdot|^{2}]=R^{2}$. For $q\geq 2$, if $\sigma^{2}\gtrsim qR\,\lVert\operatorname{cov}\pi\rVert^{1/2}\log^{1/2}d$, then $\mathcal{R}_{q}(\pi\gamma_{\sigma^{2}}\mathbin{\|}\pi)=O(1)$.
@@
By [Mil09, §5] or [CG20, Theorem 3.13], if two log-concave probability measures $\nu_{1}$ and $\nu_{2}$ satisfy $\lVert\nu_{1}-\nu_{2}\rVert_{\mathsf{TV}}\leq\varepsilon$ for $\varepsilon\in(0,1)$, then
@@
Since $\pi_{t}\to\pi$ in $\mathsf{TV}$-distance (by Scheffé), and $\pi_{t}$ is log-concave (by Prékopa–Leindler), we have $\limsup_{t\downarrow 0}C_{\mathsf{PI}}(\pi_{t})\lesssim C_{\mathsf{PI}}(\pi)$. Next, emulating the proof of Lemma 4.1 with $\alpha=\infty$, (4.3) yields
@@
Covariance of strongly log-concave distributions.
@@
Query complexity of Rényi-warmness generation.
Now that all pieces are ready, we show that an $O(1)$-warm start in $\mathcal{R}_{q}$ can be generated using $\widetilde{O}(qd^{2}R^{3/2}\Lambda^{1/4})$ membership queries for $q\geq 2$, where $B_{1}(0)\subset\mathcal{K}$, $\Lambda=\lVert\operatorname{cov}\pi\rVert$, and $R^{2}=\mathbb{E}_{\pi}[|\cdot|^{2}]$.
@@
The initialization by rejection sampling requires $O(1)$ queries in expectation due to
@@
In the annealing phase, suppose that we are given $\bar{\mu}_{i}$ such that $\mathcal{R}_{3}(\bar{\mu}_{i}\mathbin{\|}\mu_{i})=O(1)$. By a weak triangle inequality for Rényi divergence (with $\lambda=2/3$),
$\mathcal{R}_{2}(\bar{\mu}_{i}\mathbin{\|}\mu_{i+1})\leq\mathcal{R}_{3}(\bar{\mu}_{i}\mathbin{\|}\mu_{i})+\mathcal{R}_{4}(\mu_{i}\mathbin{\|}\mu_{i+1})\leq O(1)+2=O(1)\,,$
@@
where the last inequality follows from Lemma 4.1. By Theorem 1.4, from an $O(1)$-warmness in $\mathcal{R}_{2}$-divergence, $\mathsf{PS}_{\textup{Gauss}}$ needs $O(d^{2}\sigma_{i}^{2})$ queries to sample $X_{i+1}^{*}\sim\bar{\mu}_{i+1}$ such that $\mathcal{R}_{3}(\bar{\mu}_{i+1}\mathbin{\|}\mu_{i+1})=O(1)$. Also, doubling of initial $\sigma_{i}^{2}$ requires $O(R/\sigma_{i})$ phases, so each doubling needs $O(d^{2}\sigma_{i}R)$ queries in total. Since there are at most $O(\log qdR\Lambda^{1/2})$ doublings, and $\sigma_{i}^{2}\leq\sigma_{\text{last}}^{2}:=O(qR\Lambda^{1/2}\log^{1/2}d)$, we can obtain a sample $X^{*}$ such that $\mathcal{R}_{3}(\bar{\mu}_{\text{last}}\mathbin{\|}\mu_{\text{last}})=O(1)$, using
@@
$O\bigl{(}q^{1/2}d^{2}R^{3/2}\Lambda^{1/4}\log^{1/4}d\times\log qdR\Lambda^{1/2}\bigr{)}=\widetilde{O}(q^{1/2}d^{2}R^{3/2}\Lambda^{1/4})\quad\text{queries}\,.$
@@
By hypercontractivity of $\mathsf{PS}$ (Corollary 1.3), iterating $\mathsf{PS}_{\textup{Gauss}}$ started from $X^{*}$ towards $\mu_{\text{last}}$ $O(d^{2}\sigma_{\text{last}}^{2}\log_{2}q)$ times leads to a new sample $X^{\prime}$ such that $\mathcal{R}_{2q}(\operatorname{law}X^{\prime}\mathbin{\|}\mu_{\text{last}})=O(1)$. Hence, the entire annealing has used
$\widetilde{O}\bigl{(}q^{1/2}d^{2}R^{3/2}\Lambda^{1/4}+qd^{2}R\Lambda^{1/2}\bigr{)}\leq\widetilde{O}\bigl{(}qd^{2}R^{3/2}\Lambda^{1/4}\bigr{)}\quad\text{queries (due to }R^{2}\geq\Lambda\text{)}\,.$
@@
Lastly, we conclude the proof by noting the early-stopping result (Lemma 4.2) and the weak triangle inequality, which results in $\mathcal{R}_{q}(\operatorname{law}X^{\prime}\mathbin{\|}\pi)=O(1)$. ∎
5 Log-concave sampling under an evaluation oracle

### v2

@@ 4.1 Proximal sampler for truncated Gaussian
Combining all these together, we can prove Theorem 1.4 on the query complexity of $\mathsf{PS}_{\textup{Gauss}}$.
@@
Take $\eta=\varepsilon<0.01$, and restart the algorithm upon failure. To invoke the proposition above, $q_{0}$ should satisfy $q_{0}\geq 2\vee O(\log N)$ for $N=\widetilde{O}(d^{2}\sigma^{2}\log\frac{q}{\varepsilon}\log\frac{1}{\varepsilon})$. To ensure $q_{0}\geq\log\widetilde{O}(d^{2}\sigma^{2}\log\frac{q}{\varepsilon}\log\frac{1}{\varepsilon})(\geq\log N)$, it suffices to enforce
$q_{0}\geq 2\vee\widetilde{O}\bigl(\log d^{2}\sigma^{2}+\log\log\frac{q}{\varepsilon}\bigr)\,$,
so we will set $q_{0}$ to this as a required base order. We then restrict to $q\geq 2\vee\widetilde{O}(\log d^{2}\sigma^{2}+\log\log\frac{1}{\varepsilon})$ so that $q\geq q_{0}$ holds. Therefore, by Proposition 4.1, the restart version of $\mathsf{PS}_{\textup{Gauss}}$ uses $\widetilde{O}(d^{2}\sigma^{2}\log\frac{q}{\varepsilon}\log^{5}\frac{1}{\varepsilon})$ total queries in expectation, and its output is distributed as $\hat{\mu}_{N}^{X}$ satisfying $\mathcal{R}_{q}(\hat{\mu}_{N}^{X}\mathbin{\|}\pi^{X})\leq 3\varepsilon$. Retaking $\varepsilon\leftarrow\varepsilon/3$, we complete the proof. ∎
@@ 4.2 Functional inequalities to closeness and early-stopping
To avoid any regularity issues, we work with mollifications of logconcave distributions. As detailed in [EVA10, §C.4], the standard mollifier $\eta\in C^{\infty}(\mathbb{R}^{d})$ is defined as
[…]
@@ A Rényi version of LSI was established in [VW23, Lemma 5] that for $q>1$ and smooth $\mu/\nu$,
Let $\pi$ be a logconcave probability measure over $\mathbb{R}^{d}$ with $\mathbb{E}_{\pi}[|\cdot|^{2}]=R^{2}$. For $\alpha,\sigma^{2}>0$ and $q>1$, probability measures $\pi\gamma_{\sigma^{2}}$ and $\pi\gamma_{\sigma^{2}(1+\alpha)}$ satisfy
@@
$\frac{\int|\cdot|^{2}\,\gamma_{\delta}\,\mathrm{d}\rho\,}{\int\gamma_{\delta}\,\mathrm{d}\rho}\leq\frac{\int|\cdot|^{2}\,\mathrm{d}\rho}{\int\mathrm{d}\rho}=\mathbb{E}_{\rho}[|\cdot|^{2}]\,.$ (4.5)
Observe that $\operatorname{cov}_{\rho}(f,g)=\frac{1}{2}\,\mathbb{E}_{(X,Y)\sim\rho^{\otimes 2}}[\{f(X)-f(Y)\}\,\{g(X)-g(Y)\}]$. As $f(\cdot)=|\cdot|^{2}$ increases in $|\cdot|$ while $g=\gamma_{\delta}$ decreases, $0\geq\operatorname{cov}_{\rho}(f,g)=\mathbb{E}_{\rho}[fg]-\mathbb{E}_{\rho}f\,\mathbb{E}_{\rho}g$, which proves the claim. Using this fact with $\rho=\pi$, we complete the proof of the main claim. ∎
@@
Let $\pi$ be a logconcave probability measure over $\mathbb{R}^{d}$ with $\mathbb{E}_{\pi}[|\cdot|^{2}]=R^{2}$. For $q\geq 2$, if $\sigma^{2}\gtrsim qR\,\lVert\operatorname{cov}\pi\rVert^{1/2}\log^{1/2}d$, then $\mathcal{R}_{q}(\pi\gamma_{\sigma^{2}}\mathbin{\|}\pi)=O(1)$.
@@
By [MIL09, §5] or [CG20, Theorem 3.13], if two logconcave probability measures $\nu_{1}$ and $\nu_{2}$ satisfy $\lVert\nu_{1}-\nu_{2}\rVert_{\mathsf{TV}}\leq\varepsilon$ for $\varepsilon\in(0,1)$, then
@@
Since $\pi_{t}\to\pi$ in $\mathsf{TV}$-distance (by Scheffé), and $\pi_{t}$ is logconcave (by Prékopa–Leindler), we have $\limsup_{t\downarrow 0}C_{\mathsf{PI}}(\pi_{t})\lesssim C_{\mathsf{PI}}(\pi)$. Next, emulating the proof of Lemma 4.2 with $\alpha=\infty$, (4.4) yields
@@
Covariance of strongly logconcave distributions.
@@
4.3 Rényi-warmness generation
Now that all pieces are ready, we show that an $O(1)$-warm start in $\mathcal{R}_{q}$ (equivalently, $L^{q}=O(1)$) can be obtained by using $\widetilde{O}(qd^{2}R^{3/2}\Lambda^{1/4})$ membership queries, where $B_{1}(0)\subset\mathcal{K}$, $\Lambda=\lVert\operatorname{cov}\pi\rVert$, and $R^{2}=\mathbb{E}_{\pi}[|\cdot|^{2}]$.
Algorithm.
Given $\varepsilon_{\textup{unif}}=\eta_{\textup{unif}}\in(0,1/10)$ and $q\geq 2$, we pick a “base” order $q_{0}(\leq q)$ that will be preserved through annealing. To get a failure-free algorithm, we restart the following algorithm from scratch if $\mathsf{PS}_{\textup{Gauss}}$ declares failure in the middle: for $m$ the number of total phases,
  1. 1.
    Sample from $\mu_{1}=\pi\gamma_{d^{-1}}$ by rejection sampling with proposal $\gamma_{d^{-1}}$, and set $\bar{\mu}_{1}:=\mu_{1}$.
  2. 2.
    While $\sigma_{i}^{2}\lesssim qR\Lambda^{1/2}\log^{1/2}d$, move from $\bar{\mu}_{i}$ to $\bar{\mu}_{i+1}$ by $\mathsf{PS}_{\textup{Gauss}}$ such that $\mathcal{R}_{q_{0}}(\bar{\mu}_{i+1}\mathbin{\|}\mu_{i+2})\leq 1$, where
    $\sigma_{i+1}^{2}=\sigma_{i}^{2}\,\bigl(1+\frac{\sigma_{i}}{q_{0}R}\bigr)\,.$
    1. (a)
      For $i<m$, run $\mathsf{PS}_{\textup{Gauss}}:\mathcal{R}_{q_{0}}\to\mathcal{R}_{2q_{0}}$ with $N_{i}=\widetilde{O}(d^{2}\sigma_{i}^{2}\log q_{0}\log m)$, where we set $\varepsilon_{i}=1/10$ and $\eta_{i}=(10m)^{-1}$.
    2. (b)
      For $i=m$, run $\mathsf{PS}_{\textup{Gauss}}:\mathcal{R}_{q_{0}}\to\mathcal{R}_{q}$ with $N_{i}=\widetilde{O}(qd^{2}R\Lambda^{1/2}\log m)$, where we set $\varepsilon_{i}=1/10$ and $\eta_{i}=(10m)^{-1}$.
Note that $N_{i}$ is chosen to ensure the target accuracies of $\mathsf{PS}_{\textup{Gauss}}$ in each phase (see (4.1)). As seen shortly, given $q\geq\widetilde{O}(1)$, we can take $q_{0}=O(\log q)$ so that the annealing has “linear” dependence on qq. [^2] If we naïvely attempt to preserve $q$-warmness, then the last doubling of $\sigma_{i}^{2}$ incurs (roughly) $qd^{2}\sigma_{m}R\lesssim q^{3/2}d^{2}R^{3/2}\Lambda^{1/4}$ queries (i.e., $q^{3/2}$-dependence).
Choice of parameters.
As $m=\widetilde{O}(q_{0}d^{1/2}R\log q)$, we have $N_{i}\leq\widetilde{O}(qd^{2}R\Lambda^{1/2})=:N_{\max}(q)$. Also, $N_{\textup{unif}}:=\widetilde{O}(qd^{2}\Lambda\log\frac{1}{\eta_{\textup{unif}}}\log\frac{1}{\varepsilon_{\textup{unif}}})$. Recall that $q_{0}$ and $q$ should satisfy sufficient conditions to invoke the established sampling guarantees of $\mathsf{PS}_{\textup{Gauss}}$ and $\mathsf{PS}_{\textup{unif}}$ (Proposition 2.8 and 4.1). Namely, $q\geq 2\vee O(\log\frac{N_{\textup{unif}}}{\eta_{\textup{unif}}})$, $q_{0}\geq 2\vee O(\log mN_{\max})$, and $q\geq q_{0}$. The second condition is implied by $q_{0}\gtrsim\log\widetilde{O}(q_{0}qd^{5/2}R^{2}\Lambda^{1/2}\log q)$, so we will set $q_{0}=\widetilde{O}(\log qdR\Lambda^{1/4})$. To ensure $q\geq q_{0}$, we will restrict $q\geq\widetilde{O}(\log dR\Lambda^{1/4})$. Now, using (2.3) for the first condition, we will assume that
$q\geq 2\vee\widetilde{O}(\log dR\Lambda^{1/4})\vee\widetilde{O}\bigl(\log\frac{d^{2}\Lambda}{\eta_{\textup{unif}}}+\log\log\frac{1}{\varepsilon_{\textup{unif}}}\bigr)\,,\quad q_{0}=\widetilde{O}(\log qdR\Lambda^{1/4})\,.$ (4.7)
By abuse of notation, we sometime write $q\geq\widetilde{O}(1)$ and $q_{0}=\widetilde{O}(\log q)$. We note that $\eta_{i}$ is chosen to bound by $1/10$ the probability that the failure condition is met until the termination of the annealing.
High-level idea of analysis.
We consider an annealing process across a sequence of target distributions $\{\mu_{i}\}_{i=1}^{m}$, where each phase invokes the capped $\mathsf{PS}_{\textup{Gauss}}$ designed to move from (approximately) $\mu_{i}$ to (approximately) $\mu_{i+1}$. Each phase may declare failure due to a capped inner rejection sampler.
Let $S^{(i)}$ be the event that phase $i$ succeeds (no failure of $\mathsf{PS}_{\textup{Gauss}}$ within that phase), and let $S:=\cap_{i=1}^{m-1}S^{(i)}$ be the global success event. Let $C^{(i)}\geq 0$ be the query complexity incurred by phase $i$, and define the one-attempt total cost
$C:=\sum_{i=1}^{m-1}\mathds{1}\bigl\{\bigcap_{j=1}^{i-1}S^{(j)}\bigr\}\,C^{(i)}\,.$
Clearly, $\mathbb{E}C\leq\sum_{i=1}^{m-1}\mathbb{E}C^{(i)}$, so it suffices to upper bound $\mathbb{E}C^{(i)}$. Similarly, letting $F^{(i)}$ denote the event that phase $i$ fails and $q^{(i)}$ denote its cap-hit probability inside phase $i$, we have
$\mathbb{P}(F^{(i)})=\mathbb{E}\bigl[\mathds{1}\bigl\{\cap_{j=1}^{i-1}S^{(j)}\bigr\}\,q^{(i)}\bigr]\leq\mathbb{E}q^{(i)}\,,\qquad\mathbb{P}(S^{c})\leq\sum_{i=1}^{m-1}\mathbb{E}q^{(i)}\,$,
which allows us to bound the overall failure probability via unconditional phase-wise estimates (under the uncapped intermediate laws used in the analysis).
Similar to the analysis of the restarting version of $\mathsf{PS}_{\textup{unif}}$, a restart-until-success annealing repeats independent full attempts (including resampling the initial input from $\mu_{1}$) until $S$ occurs. If $\mathbb{P}(S)\geq 1-\eta$, then the expected number of attempts is at most $(1-\eta)^{-1}$ and
$\mathbb{E}C_{\text{restart}}=\frac{\mathbb{E}C}{\mathbb{P}(S)}\leq\frac{\mathbb{E}C}{1-\eta}\,.$
In summary, we first establish (phase-wise) bounds on unconditional failure probabilities and unconditional expected costs, combine them to control $\mathbb{P}(S)$ and $\mathbb{E}C$ for one attempt, and then translate these into guarantees for the returned-sample distribution $\nu$ and for the restart wrapper.
@@ 4.3 Rényi-warmness generation
Pick $q$ and $q_{0}$ as in (4.7). The initialization by rejection sampling requires $O(1)$ queries in expectation due to
@@
In the annealing phase, suppose that we are given $\bar{\mu}_{i}$ such that $\mathcal{R}_{2q_{0}}(\bar{\mu}_{i}\mathbin{\|}\mu_{i})\leq\frac{1}{10}$. By a weak triangle inequality for Rényi divergence (with $\lambda=2/3$),
$\mathcal{R}_{q_{0}}(\bar{\mu}_{i}\mathbin{\|}\mu_{i+1})\leq 1.1\mathcal{R}_{2q_{0}}(\bar{\mu}_{i}\mathbin{\|}\mu_{i})+\mathcal{R}_{2q_{0}}(\mu_{i}\mathbin{\|}\mu_{i+1})\leq\frac{1.1}{10}+\frac{1}{2}\leq 1\,$,
@@
where the last inequality follows from Lemma 4.2. By Theorem 1.4, from an $O(1)$-warmness in $\mathcal{R}_{q_{0}}$-divergence (precisely, $\mathcal{R}_{q_{0}}(\bar{\mu}_{i}\mathbin{\|}\mu_{i+1})\leq 1$), $\mathsf{PS}_{\textup{Gauss}}$ needs $\widetilde{O}(d^{2}\sigma_{i}^{2}\log^{2}q_{0})$ queries to sample $X_{i+1}^{*}\sim\bar{\mu}_{i+1}$ such that $\mathcal{R}_{2q_{0}}(\bar{\mu}_{i+1}\mathbin{\|}\mu_{i+1})\leq\frac{1}{10}$. Also, doubling of given $\sigma_{i}^{2}$ requires $O(q_{0}R/\sigma_{i})$ phases, so each doubling needs $\widetilde{O}(q_{0}d^{2}\sigma_{i}R)$ queries in total. Since there are at most $O(\log qdR\Lambda^{1/2})$ doublings, and $\sigma_{i}^{2}\leq\sigma_{m}^{2}\asymp qR\Lambda^{1/2}\log^{1/2}d$, we can obtain a sample $X^{*}\sim\bar{\mu}_{m-1}$ such that $\mathcal{R}_{2q_{0}}(\bar{\mu}_{m-1}\mathbin{\|}\mu_{m-1})\leq\frac{1}{10}$, using
@@
$O\bigl(q_{0}q^{1/2}d^{2}R^{3/2}\Lambda^{1/4}\log^{1/4}d\times\log qdR\Lambda^{1/2}\bigr)=\widetilde{O}(q^{1/2}d^{2}R^{3/2}\Lambda^{1/4})\quad\text{queries}\,.$
@@
By hypercontractivity of $\mathsf{PS}$ (Corollary 1.3), iterating $\mathsf{PS}_{\textup{Gauss}}$ started from $X^{*}$ towards $\mu_{m}$ $\widetilde{O}(d^{2}\sigma_{m}^{2}\log q)$ times leads to a new sample $X^{\prime}$ such that $\mathcal{R}_{2q}(\operatorname{law}X^{\prime}\mathbin{\|}\mu_{m})\leq\frac{1}{10}$. Hence, the entire annealing has used
$\widetilde{O}\bigl(q^{1/2}d^{2}R^{3/2}\Lambda^{1/4}+qd^{2}R\Lambda^{1/2}\bigr)\leq\widetilde{O}(qd^{2}R^{3/2}\Lambda^{1/4})\quad\text{queries (due to }R^{2}\geq\Lambda\text{)}\,.$
@@
Note that the early-stopping result (Lemma 4.3) and the weak triangle inequality imply $\mathcal{R}_{q}(\operatorname{law}X^{\prime}\mathbin{\|}\pi)\leq 1$ as desired.
Lastly, restarting the annealing (in case of failure) increases the total query complexity multiplicatively by $10/9$, and the conditional law on success has an extra bias at most $1$ in $\mathcal{R}_{\infty}$. Hence, by $\mathcal{R}_{q}(\operatorname{law}(X^{\prime}\mid\text{success})\mathbin{\|}\pi)\leq 2$. ∎
We now establish the query complexity of convex-body sampling from scratch.
Proof of Corollary 1.6.
We set $\eta_{\textup{unif}}=\varepsilon_{\textup{unif}}=\varepsilon<1/100$ and consider $q\geq\widetilde{O}(1)$. By Theorem 1.1, the expected number of restarts of $\mathsf{PS}_{\textup{unif}}$ is at most $O(1)$, so the warm-start generation will be repeated at most $O(1)$ times. Combining the complexity results in Theorem 1.5 and 1.1, the total complexity of logconcave sampling from scratch is
$\widetilde{O}\bigl(qd^{2}R^{3/2}\Lambda^{1/4}+qd^{2}\Lambda\log^{6}\frac{1}{\varepsilon}\bigr)\,.$
When $1\leq q\leq 2\vee\widetilde{O}(\log\frac{dR}{\varepsilon})=:q_{*}$, we generate the $\mathcal{R}_{q_{*}}$-warmness and use the monotonicity of $\mathcal{R}_{q}$. ∎
5 Logconcave sampling under an evaluation oracle